# Low Variance Off-policy Evaluation with State-based Importance Sampling

### v4

@@ 3.3 Applying off-policy evaluation
Off-policy evaluation has two primary uses. First, when few trajectories are available and the designer wants to make a decision on whether or not to use the evaluation policy instead of a given expert policy; primary examples include medical treatment [2] and digital marketing [20], which would incur significant costs and safety risks when the evaluation policy’s performance is not estimated correctly. Second, off-policy evaluation can be used to optimise the policy in what is called off-policy improvement. Important topics in this area are how to ensure monotonic improvement and how to provide confidence estimates on improvement. Off-policy improvement is based on batch algorithms since these take into account the trajectories sampled from the behaviour policy. High-confidence policy improvement (HCOPE) [24] uses confidence estimates on the evaluation [23] to ensure with high probability that the new policy improves on the old. Policy optimisation via importance sampling (POIS) [10] optimises surrogate objective that effectively captures the trade-off between the estimated improvement and the uncertainty due to importance sampling using the 2-Renyi divergence of the evaluation policy and the behaviour policy. POIS benefits from using multiple importance sampling and per-decision importance sampling [11] so a possible integration with state-based methods may also be considered in future work. Uniform OPE [27] provides stronger guarantees for off-policy evaluation by simultaneously evaluating all policies in a policy class, thereby integrating naturally with policy improvement.
4 State-based Importance Sampling
State-based importance sampling (SIS) mitigates the above issue by constructing an estimator $\hat{G}_{SIS}$ that selectively drops the action probability ratios of selected states from the product to compute the importance weight, such that 1) on expectation, the dropped term is equal to 1; and 2) the dropped term does not covary with the estimator $\hat{G}_{SIS}$. Additionally, we note that the variance upper bound is lowered if the remaining action probability ratios have a low upper bound, although this does not affect the estimator’s bias. The theory behind this is outlined below.
@@ 4 State-based Importance Sampling
where $[H]=\{1,2,\dots,H\}$. Then, following the covariance testing approach by Guo et al. (2017) [5], the expected return can be rewritten as
$\mathscr{G}$ $=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[G\left(\prod_{t=1}^{H}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right)\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[GAB\right]$
$=\mathbb{E}_{\pi_{b},\mathcal{T}}\left[A\right]\mathbb{E}_{\pi_{b},\mathcal{T}}\left[BG\right]+\text{Cov}(A,BG)\,,$
where the last step follows from $\text{Cov}(X,Y)=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]$. The above implies that when provided with a sufficient number of samples from the random variable $BG$, one can accurately estimate $\mathscr{G}$ based on samples of $BG$ whenever $\mathbb{E}_{\pi_{b}}\left[A\right]=1$ and $\text{Cov}(A,BG)=0$.
Condition 1: $\mathbb{E}_{\pi_{b}}\left[A\right]=1$ This condition holds if
$\mathbb{E}_{\pi_{b},\mathcal{T}}\left[\prod_{t\in[H]:s_{t}\in\mathcal{S}^{A}}\frac{\pi_{e}(a_{t}|s_{t})}{\pi_{b}(a_{t}|s_{t})}\right]=1\,.$
The transition dynamics model $\mathcal{T}$ of the MDP is not known and therefore the frequency of states in $\mathcal{S}^{A}$ occurring in these trajectories is unknown and may vary considerably. However, the random variable $A$ can be seen as the importance weight of an evaluation policy $\pi_{e}$ that is equal to $\pi_{e}$ for $s\in\mathcal{S}^{A}$ and equal to $\pi_{b}$ for $s\in\mathcal{S}^{B}$. Therefore, $A$ can be seen as an importance weight of the entire trajectory under a different evaluation policy $\pi_{e}^{A}$, which we will call the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$. Proof is given in Appendix A.
Condition 2: $\text{Cov}(A,BG)=0$ This condition holds if $BG$ does not grow bigger (resp. smaller) if $A$ increases (resp. decreases). Condition 2 can be verified according to the algorithm proposed in Section 5; while the algorithm verifies it explicitly, it may also be done implicitly due to the contribution to the MSE.
Condition 3: the states $s\in\mathcal{S}^{B}$ have low maximal action probability ratio.
Provided that for some $\epsilon>0$, $\pi_{b}(s,a)\geq\epsilon$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and that $\pi_{e}\leq 1$, a general upper bound to the action probability ratio is $\rho_{\text{max}}=1/\epsilon^{H}$. Therefore, this condition is satisfied in general when a uniform random policy is chosen for the behaviour policy $\pi_{b}$, since these have $\epsilon=1/|\mathcal{A}|$. This may not always desirable, e.g. when requiring a safe expert policy, but for variance reduction in SIS this can be an important factor for improving worst-case bounds. Note that Condition 3 has a beneficial effect on variance reduction but has no influence on the bias. A minimalistic interpretation of Condition 3 would to have a soft behaviour policy (i.e. with no zero-probability actions) such that no importance ratios can be infinite, which is a typical assumption in importance sampling approaches to off-policy evaluation.
Under the above conditions, the SIS estimator is constructed for a given dropped state set $\mathcal{S}^{A}$ as
@@
Rather than exponential dependence on $H$, this variance upper bound of $\hat{G}_{SIS}$ is exponentially dependent on the maximal number of occurrences of states in $\mathcal{S}^{B}$. Denoting the subtrajectory $\tau_{B}:=\{t\in[H]:s_{t}^{(i)}\in\mathcal{S}^{B}\}$ and $M_{B}:=\max_{\tau\in\tau_{B}}|\tau|$, we have
@@
This result implies that the performance of the SIS estimator will be significantly improved if $M_{B}$ is sufficiently small, i.e. if state dropping is a frequent occurrence.
The underlying principle behind covariance testing [5] is that dropping a term $A$ from the importance-weighted return $ABG$ for an estimate $BG$ introduces a bias that is equivalent to the covariance $\text{Cov}(A,BG)$. Consequently, the MSE is given by
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\text{Var}(\hat{G}_{SIS}(\mathcal{S}^{A}))+\text{Cov}(A,BG)^{2}\,.$ (6)
This leads to the definition of negligibility below.
@@
$(\pi_{e},\pi_{b},\epsilon)$-decomposable MDP and $\epsilon$-negligible state set. Let $\epsilon>0$, $\mathcal{M}$ be an MDP $\langle\mathcal{S},\mathcal{A},r,\mathcal{T}\rangle$, and let $\pi_{e}$ and $\pi_{b}$ two policies in $M$. Then $\mathcal{M}$ is $(\pi_{e},\pi_{b},\epsilon)$-decomposable iff there exists a set $\mathcal{S}^{A}$ such that $\text{Cov}(A,BG)<\epsilon$. Any such set $\mathcal{S}^{A}$ is called an $\epsilon$-negligible state set for the off-policy evaluation problem $\langle\mathcal{M},\pi_{e},\pi_{b}\rangle$.
When $\mathcal{S}^{A}$ is an $\epsilon$-negligible state set for an off-policy evaluation problem, this implies via Equation 56 that
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))\leq\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}+\epsilon^{2}\,.$ (7)
5 Optimising the dropped state set
We propose two search algorithms to optimise the dropped state set, one based on covariance testing and one based on state-action values (a.k.a. Q-values). Each search algorithm comes with its own identification of negligible states, i.e. those that are believed to be allowed to drop with only a negligible increase in bias. Inaccurate identification of negligible state sets can increase the MSE significantly. However, as illustrated in Appendix B, if many states can be dropped then despite this additional bias, the performance of SIS is often a significant improvement over IS.
5.1 Covariance testing
To identify $\epsilon$-negligible state sets, one can take the approach taken in incremental importance sampling (INCRIS) [5], which is to compute an empirical estimate of the mean squared error (MSE) based on the bias-variance decomposition of the MSE:
$\widehat{\text{MSE}}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\widehat{\text{Var}}(\hat{G}_{SIS}(\mathcal{S}^{A}))+\widehat{\text{Cov}}(A,BG)^{2}\,,$ (8)
where $\widehat{\text{Cov}}(A,BG)$ is the empirical covariance estimate and $\widehat{\text{Var}}(A)$ is the empirical variance estimate. In SIS, the state set $\mathcal{S}^{A}$ to be dropped from the importance weight computation may require checking many subsets of states within $\mathcal{S}$. In this paper, we perform experiments in a limited state space which allows performing an exhaustive search over many subsets – although rather than the full power set, we limit the search to sets with cardinalities of at most 2. The algorithm defines the best state set incrementally by defining a state set $\texttt{best}\leftarrow\mathcal{S}^{A}$ whenever either $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})$ or $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})*(1+\epsilon_{s})\land|\mathcal{S}^{A}|>|\texttt{best}|$ based on a user-defined $\epsilon_{s}>0$. Before doing so the state set must satisfy Condition 2, which is verified by checking $\epsilon$-negligibility using the estimated covariance by $\widehat{\text{Cov}}(A,BG)<\epsilon$.
5.2 Q-value based identification
Another option we explore as a search algorithm to identify negligible states is to keep track of a state-action value table which indicates the expected cumulative reward from a particular state-action pair. The reasoning is that, if the state’s action probability ratio is to be dropped, it should have state-action values that are highly similar. To do this end, a model $M$ of the MDP is formulated based on the same trajectories that are given for off-policy evaluation. The model estimates, based on the observed state visitation frequencies and reward intake, for each $s\in\mathcal{S}$ the initial distribution $\hat{P}_{1}(s)$ and for each $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ the transition probability $\hat{\mathcal{T}}(s,a,s^{\prime})$ as well as the estimated average reward $\hat{r}(s,a)$. Using the model, one can then compute time-specific Q-values via stochastic dynamic programming:
where :::: indicates vectorisation.
Intuitively, the action probability ratios for a state for which actions have similar Q-values can be ignored for the purposes of computing $\hat{G}$ since the same cumulative value would be obtained regardless of the action. This leads to the Q-value based algorithm, which identifies a state $s\in\mathcal{S}$ as negligible if for all $t\in[H]$, $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ for all $a,a^{\prime}\in\mathcal{A}$.
More formally, a simple proof can show that using this criterion leads to an SIS estimator with limited absolute bias depending on the user-defined $\epsilon$, the maximal number of dropped weights, and the model approximation error $\epsilon_{2}$ (see Appendix C). Consistent with the analysis in Appendix B, the MSE of SIS is $O\left(\exp(M_{B})\right)$ with both state identification methods.
The Q-value based identification algorithm has two main advantages over the covariance testing framework. First, the algorithm is computationally inexpensive and makes it easy to include a large subset of the state space as negligible states. Second, the algorithm is easy to interpret and independent of the estimator implying it is easily used on different importance sampling frameworks such as per-decision importance sampling and incremental importance sampling.
6 State-based importance sampling variants
While so far we have focused on ordinary importance sampling, it is easy to extend SIS to use more advanced OPE estimators as its base-estimator.
First, methods such as PDIS and INCRIS are even the various weighted counterparts of these algorithms are easily extended by analogously setting the action probability ratio to one for negligible states.
Second, doubly robust off-policy evaluation [6, 21] can also benefit from state-based importance sampling with only minor adjustments. The model of the MDP and the time-specific Q-value used in the Q-value based search algorithm is based on the same formalisms that underly doubly robust (DR) off-policy evaluation. While traditional DR uses an importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ for a trajectory $i$ at time $t$ and $n$ total trajectories, doubly robust state-based importance sampling (DRSIS) uses the importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}^{A}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ where $\pi_{e}^{A}$ is the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$. To illustrate the simplicity of this change, consider that DR computes the temporary variable $\rho_{t}^{i}$ such that $\rho_{t}^{i}=\pi_{e}(a_{t}^{i}|s_{t}^{i})/\pi_{b}(a_{t}^{i}|s_{t}^{i})*\rho_{t-1}^{i}$ and the only modification for DRSIS is to set $\rho_{t}^{i}=\rho_{t-1}^{i}$ if $s_{t}^{i}\in\mathcal{S}^{A}$.
Third, techniques that rely on importance sampling over state visitation frequencies such as SDRE and MIS can also benefit from a state-based formulation. For example, the stationary density ratio estimator (Eq. 1) can be turned into a state-based formulation according to
where $\hat{d}^{\pi_{e}^{A}}(s_{t}^{(i)})$ is equal to $\hat{d}^{\pi_{e}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{B}$ and $\hat{d}^{\pi_{b}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{A}$. Such techniques are polynomial in $H$ due to replacing the product over action probability ratios by a sum of state visitation density ratios. As such the variance reduction due to reducing polynomial factors is smaller than variance reductions due to reducing the exponent, as is the case in traditional IS methods. However, as the experiments in Section 7 show, this can still yield significant MSE improvements and an excellent overall performance in environments with stationary distributions.
7 Experiments
Having introduced various state-based importance sampling techniques, we now turn to validating them experimentally. Since state-based techniques drop particular states from the computation, one would expect these to be advantageous on domains where there are some states for which the action taken has no or limited impact on the return. This intuition is captured by the notion of “lift states”.
Definition 2.
@@ 7 Experiments
Lift states. A lift state $s\in\mathcal{S}$ is a state for which $\mathcal{T}(s,a)=\mathcal{T}(s,a^{\prime})$ and $r(s,a)=r(s,a^{\prime})$ for all pairs of actions $a,a^{\prime}\in\mathcal{A}$. The set of lift states for a domain is denoted by $\mathcal{S}^{L}$. Similarly, an approximate lift state $s\in\mathcal{S}$ is a state for which $\mathcal{T}(s,a)\approx\mathcal{T}(s,a^{\prime})$ and $r(s,a)\approx r(s,a^{\prime})$ for all pairs of actions $a,a^{\prime}\in\mathcal{A}$.
Lift states intuitively capture the covariance condition (see Condition 2) since the action probabilities, and therefore their ratios, do not affect the expected return. Analogously, for approximate lift states, action probabilities have a negligible effect on the expected return.
The experimental validation of SIS comprises 4 domains. Guided by the above intuition of lift states, the first two domains include lift states in the form of a lift transporting the agent from one side to another. As state-based importance sampling techniques are expected to be general purpose, we also assess them on two alternative domains. The third and fourth domain are common RL benchmarks, namely the inventory management domain [16] and the taxi domain variant used in [8]. Code with all estimators and experiments is available at https://github.com/bossdm/ImportanceSampling.
The experiments validate state-based importance sampling when implemented on different variants of importance sampling:
  • ordinary importance sampling (IS);
  • per-decision importance sampling (PDIS; [15]);
  • incremental importance sampling (INCRIS; [5]);
  • doubly robust OPE (DR; [21]); and
  • stationary density ratio estimation (SDRE; [8]).
The corresponding state-based implementations are denoted with the prefix S; for instance, SIS and SPDIS denote state-based IS and state-based PDIS. State-based doubly robust OPE is denoted as DRSIS to help distinguish it more clearly from SDRE. With the exception of the first domain, which has short trajectory lengths, and the SDRE technique for which this is not applicable, the experiments use additional variance reduction using weighted importance sampling [14], a technique which divides by the sum of importance weights rather than the number of trajectories. Weighted estimators are given the prefix W; for instance, WIS denotes weighted IS.
Neglible state identification parameters are fixed for each domain. Covariance testing is applied with $\epsilon_{s}=\epsilon=0.01$ in all the domains except the taxi domain for which the covariance test is too expensive to compute. The Q-value based identification is suitable for all the domains and the parameter settings are $\epsilon=1$ in the grid world lift domains, $\epsilon=50$ in the inventory management domain, and $\epsilon=2$ in the taxi domain. The effectiveness of the negligible state identification is demonstrated on the first two domains, where the performance based on state identification methods can be compared to the performance based on the lift states of these domains. The comparison for traditional importance sampling includes SIS (Lift states), which uses the lift states of the domain (i.e., where $\mathcal{S}^{A}=\mathcal{S}^{L}$); SIS (Covariance testing), which uses the negligible states identified by covariance testing; and SIS (Q-based), which uses the negligible states identified using Q-values. Similarly, we evaluate the effectiveness of state identification for the doubly robust OPE method, which are analgously denoted as SDR (Lift states), SDR (Covariance testing), and SIS (Q-based). The experiments demonstrate the Q-based identification method to correctly identify lift states and yield higher performance, which is why by default the Q-based identification is used unless indicated by the terms in parentheses.
7.1 Deterministic lift domain
The deterministic lift domain consists of consecutive lift states organised symmetrically around the starting point and with deterministic state transition dynamics. The domain involves an agent moving on a straight line, with the state space $\mathcal{S}=\{-b,\dots,-1,0,1,\dots,b\}$ defined by the bound $b$ representing the coordinates along the line and the action space $\mathcal{A}=\{\texttt{left},\texttt{right}\}$ representing left or right movements on the line. The agent starts from a central square with coordinate $s=0$ and each action moves one step left $T(s,\texttt{left},s-1)=1$ or one step right $T(s,\texttt{left},s+1)=1$ each with probability $1$ except from lift states. The episode stops when the agent hits the bound, then the terminal reward of $r=-c$ is given for if $s=-c$ and $r=c$ if $s=c$. All non-terminal rewards are $r=-1$ to penalise the length of the path. In this domain, lift states are introduced in the states $\mathcal{S}^{L}=\{-c+2,\dots,-1,1,\dots,c-2\}$, with $s\in\mathcal{S}^{L}\cap\mathbb{N}^{-}$ deterministically moving the agent further left (i.e. $T(s,\texttt{left},s-1)=T(s,\texttt{right},s-1)=1$) and $s\in\mathcal{S}^{L}\cap\mathbb{N}^{+}$ automatically moving the agent further right (i.e. $T(s,\texttt{left},s+1)=T(s,\texttt{right},s+1)=1$). The experiments manipulate this bound $c$ in $\{3,4,5,6,7,8\}$, which leads to results for domain sizes of $|\mathcal{S}|\in\{7,9,11,13,15,17\}$ and corresponds to an increasing effective horizon.
Analysing the different negligible state set identification methods reveals that SIS (Q-based) identifies the lift states as negligible states and therefore yields the same performance as SIS (Lift states). Table 5 in Appendix C demonstrates the superior performance of SIS (Q-based) and SIS (Lift states) compared to IS, which never drops any states, and SIS (Covariance testing), which typically drops two lift states (the maximal cardinality given the computational budget).

### v5

@@ III-B Applying off-policy evaluation
Off-policy evaluation has two primary uses. First, when few trajectories are available and the designer wants to make a decision on whether or not to use the evaluation policy instead of a given expert policy; primary examples include medical treatment [6] and digital marketing [23], which would incur significant costs and safety risks when the evaluation policy’s performance is not estimated correctly. Second, off-policy evaluation can be used to optimise the policy, where key challenges include how to provide monotonic improvement and how to provide confidence estimates on improvement. For example, high-confidence policy improvement (HCOPE) [24] uses confidence estimates on the evaluation [25] to ensure with high probability that the new policy improves on the old; policy optimisation via importance sampling (POIS) [26] optimises surrogate objective that effectively captures the trade-off between the estimated improvement and the uncertainty due to importance sampling [22]; and Uniform OPE [27] simulateneously evaluates all policies in a policy class.
IV State-based Importance Sampling
State-based importance sampling (SIS) mitigates the above issue by constructing an estimator $\hat{G}_{SIS}$ that selectively drops the action probability ratios of a select state set $\mathcal{S}^{A}\subset\mathcal{S}$ from the product to compute the importance weight. If the dropped term is equal to 1 and does not covary with the estimator $\hat{G}_{SIS}$, this yields lower-variance estimates. In this section, we demonstrate the state-based formalism for ordinary importance sampling while Sec. VI shows how to apply the state-based formalism for other estimators with lower variance and mean squared error. We extend ordinary importance sampling here for its simplicity and the fact that its variance is more significantly reduced (i.e. by reducing the exponent).
@@ IV State-based Importance Sampling
where $[H]=\{1,2,\dots,H\}$.
Given a selected state-set $\mathcal{S}^{A}$ of dropped states, the state-based variant of ordinary importance sampling computes the action probability ratios only over its complement $\mathcal{S}^{B}$:
@@
Its expected return can be rewritten as
$\mathcal{G}$ $=\mathbb{E}_{\pi_{b}}\left[A\right]\mathbb{E}_{\pi_{b}}\left[BG\right]+\text{Cov}(A,BG)\,,$ (6)
since $\text{Cov}(X,Y)=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]$. As shown in the coviarance testing framework [8], any decomposition of the form (6) allows to accurately estimate $\mathcal{G}$ based on samples of $BG$ if $\mathbb{E}_{\pi_{b}}\left[A\right]\approx 1$ and $\text{Cov}(A,BG)$ is sufficiently small. For the first condition, the random variable $A$ can be seen as the importance weight under a different evaluation policy $\pi_{e}^{A}$, which we will call the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$, which is equal to $\pi_{e}$ for $s\in\mathcal{S}^{A}$ and equal to $\pi_{b}$ for $s\in\mathcal{S}^{B}$; therefore it follows that $\mathbb{E}[A]=1$ (see Appendix A). The second condition does not follow automatically, but depends on the chosen $\mathcal{S}^{A}$; the definition below specifies the criterion for $S^{A}$ to form an accurate state-based estimator.
Definition 1.
$(\pi_{e},\pi_{b},\epsilon)$-decomposable MDP and $\epsilon$-negligible state set. Let $\epsilon>0$, $\mathcal{M}$ be an MDP $\langle\mathcal{S},\mathcal{A},r,\mathcal{T}\rangle$, and let $\pi_{e}$ and $\pi_{b}$ two policies in $M$. Then $\mathcal{M}$ is $(\pi_{e},\pi_{b},\epsilon)$-decomposable iff there exists a set $\mathcal{S}^{A}$ for which $\text{Cov}(A,BG)<\epsilon$. Any such set $\mathcal{S}^{A}$ is called an $\epsilon$-negligible state set for the off-policy evaluation problem $\langle\mathcal{M},\pi_{e},\pi_{b}\rangle$.
Applying bias-variance decomposition, it follows that
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\text{Cov}(A,BG)^{2}+\text{Var}(\hat{G}_{SIS}(\mathcal{S}^{A}))\,.$ (7)
Denoting the subtrajectory $\tau_{B}:=\{t\in[H]:s_{t}^{(i)}\in\mathcal{S}^{B}\}$ and $M_{B}:=\max_{\tau\in\tau_{B}}|\tau|$, we have
@@
The resulting variance upper bound of $\hat{G}_{SIS}$ is exponential in the maximal number of occurrences of states in $\mathcal{S}^{B}$ rather than in the full trajectory length $H$. Consequently, the SIS estimator will significantly improve on IS if many visited states can be dropped from the computation of the importance weight. When $\mathcal{S}^{A}$ is an $\epsilon$-negligible state set, this implies via (7) and (8) that
$\text{MSE}(\hat{G}_{SIS}(\mathcal{S}^{A}))\leq\epsilon^{2}+\frac{1}{4n}\left(Hr_{\text{max}}\rho_{\text{max}}^{M_{B}}\right)^{2}\,.$ (9)
V Optimising the dropped state set
We propose two search algorithms to optimise the dropped state set, one based on covariance testing and one based on state-action values (a.k.a. Q-values). Additional derivations are presented in Appendix B showing that even with some additional bias, the performance of a state-based variant of ordinary importance sampling (SIS) is still a significant improvement over ordinary importance sampling.
V-A Covariance testing
Following the covariance testing approach [8], this algorithm applies a bias-variance decomposition based on empirical estimates of the covariance and variance, and compares the MSEs of different parameter choices. In our case, the parameter choice is the state set $S^{A}$ for a decomposition
$\widehat{\text{MSE}}(\hat{G}_{SIS}(\mathcal{S}^{A}))=\widehat{\text{Cov}}(A,BG)^{2}+\widehat{\text{Var}}(\hat{G}_{SIS}(\mathcal{S}^{A}))\,,$ (10)
and the algorithm encourages larger negligible states sets using the criterion $\texttt{best}\leftarrow\mathcal{S}^{A}$ if and only if $\widehat{\text{Cov}}(A,BG)<\epsilon$ and $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})$ or $\widehat{\text{MSE}}(\mathcal{S}^{A})<\widehat{\text{MSE}}(\texttt{best})*(1+\epsilon_{s})\land|\mathcal{S}^{A}|>|\texttt{best}|$ based on user-defined parameters $\epsilon_{s}>0$ and $\epsilon>0$. In this paper, we use exhaustive search over subsets while it is also possible to form heuristic search algorithms.
V-B Q-value based identification
In the Q-value based identification, a model of the MDP is formulated based on the trajectories from $\pi_{b}$ to compute time-specific Q-values for $\pi_{e}$ via stochastic dynamic programming:
where :::: indicates vectorisation. The algorithm identifies a state $s\in\mathcal{S}$ as negligible if for all $t\in[H]$, $|\hat{Q}_{t}(s,a)-\hat{Q}_{t}(s,a^{\prime})|<\epsilon$ for all $a,a^{\prime}\in\mathcal{A}$. This criterion leads to an SIS estimator with limited absolute bias depending on the user-defined $\epsilon$, the maximal number of dropped weights, and the model approximation error $\epsilon_{2}$ (see Appendix C). Consistent with the analysis in Appendix B, the MSE of SIS is $\mathcal{O}\left(\exp(M_{B})\right)$ with both state identification methods. The Q-value based identification algorithm is computationally inexpensive and can include a large subset of the state space as negligible states. Moreover, it easy to interpret and independent of the estimator making it applicable to a variety of estimators.
TABLE I: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the deterministic lift domain based on 50 independent runs of 1,000 episodes for each domain size. The best and second best estimators are highlighted in bold underlined and bold, respectively, for each domain size. All SIS estimators shown use the Q-value based algorithm for negligible state set identification.
Domain size IS SIS PDIS SPDIS INCRIS SINCRIS SDRE SSDRE
7 0.0067 0.0019 0.0398 0.0121 4.0024 4.0007 532.7567 536.3787
9 0.0153 0.0019 0.1766 0.0215 6.1964 6.1765 577.3089 584.4657
11 0.0353 0.0029 0.7087 0.0600 8.9408 8.9225 674.0772 680.1959
13 0.0691 0.0026 1.8038 0.0688 12.5475 12.4779 652.6004 660.2387
15 0.1389 0.0033 5.3474 0.1378 16.1902 16.1558 803.3450 813.4997
17 0.3026 0.0031 15.1566 0.1303 23.7084 20.1534 758.2237 764.6283
TABLE II: Mean squared error (MSE) for off-policy evaluation estimators of the expected return in the stochastic lift domain based on 200 independent runs of 1,000 episodes for each domain size. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each domain size.
Domain size WIS WSIS WPDIS WSPDIS WINCRIS WSINCRIS SDRE SSDRE WDR WDRSIS
7 10.2238 8.4832 6.8271 5.3853 1.0958 0.2618 208.7413 209.8354 5.2572 5.0844
9 9.4737 7.7326 5.3924 4.0441 0.2973 0.6929 243.1366 245.1004 2.8067 2.7486
11 9.7285 6.2044 5.7219 3.0552 0.2002 2.3050 250.3149 251.5843 2.3016 1.9838
13 8.4436 9.4967 4.4449 3.2916 0.8958 4.8960 266.7911 268.6634 1.3459 1.3354
15 18.0453 4.2831 6.2259 1.7004 2.6550 9.0162 280.6434 282.6849 1.2258 0.8797
17 8.2474 4.8129 5.2761 1.8151 4.7157 13.0056 305.7465 307.9191 0.8385 0.4820
VI State-based importance sampling variants
While so far we have focused on ordinary importance sampling, it is easy to extend SIS to use more advanced off-policy evaluation estimators as its base-estimator.
First, methods such as PDIS and INCRIS, and even the weighted counterparts of these algorithms, are easily extended by analogously setting the action probability ratio to one for negligible states.
Second, for doubly robust off-policy evaluation [17, 10], one can use a model using the same formalism as the Q-value based identification. While traditional DR uses an importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ for a trajectory $i$ at time $t$ and $n$ total trajectories, doubly robust state-based importance sampling (DRSIS) uses the importance weight $w_{t}^{i}=\frac{1}{n}\prod_{j=1}^{t}\frac{\pi_{e}^{A}(a_{t}^{i}|s_{t}^{i})}{\pi_{b}(a_{t}^{i}|s_{t}^{i})}$ where $\pi_{e}^{A}$ is the $\mathcal{S}^{A}$-reduced evaluation policy of $\pi_{e}$.
Third, techniques that rely on importance sampling over state visitation frequencies such as SDRE and MIS can also benefit from a state-based formulation. For example, the stationary density ratio estimator (2) can be turned into a state-based formulation according to
where $\hat{d}^{\pi_{e}^{A}}(s_{t}^{(i)})$ is equal to $\hat{d}^{\pi_{e}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{B}$ and $\hat{d}^{\pi_{b}}(s_{t}^{(i)})$ for $s_{t}^{(i)}\in\mathcal{S}^{A}$. Such techniques are polynomial in $H$ due to replacing the product over action probability ratios by a sum of state visitation density ratios. As such the variance reduction due to reducing polynomial factors is smaller than variance reductions due to reducing the exponent, as is the case in traditional IS methods. However, as the experiments in Sec. VII show, this can still yield significant MSE improvements and an excellent overall performance in environments with stationary distributions.
VII Experiments
The experimental validation of state-based estimators comprises 4 domains. The first two domains include lift states in the form of a lift transporting the agent from one side to another, while the third and fourth domain are common RL benchmarks, namely the inventory management domain [28] and the taxi domain variant used in [7]. In each of the domains, we evaluate the expected return of a near-optimal policy $\pi_{e}$ based on sub-optimal policy $\pi_{b}$.
The experiments validate state-based importance sampling when implemented on IS, PDIS, INCRIS, DR, and SDRE. The corresponding state-based implementations are denoted with the prefix S; for instance, SPDIS denotes state-based PDIS. State-based doubly robust off-policy evaluation is denoted as DRSIS to distinguish it from SDRE. With the exception of the first domain, which has short trajectory lengths, and SDRE for which this is not applicable, the experiments use additional variance reduction with weighted importance sampling, which is indicated by the prefix W; for instance, WIS denotes weighted IS. Further, the effectiveness of the negligible state identification is demonstrated on the first two domains as these contain known lift states (see Appendix D). Due to its ability to identify the lift states and yield lower MSE, the Q-based identification algorithm is used throughout this section. Additional experimental details are provided in Appendix E and source code is available at https://github.com/bossdm/ImportanceSampling.
TABLE III: Normalised mean squared error (MSE) for off-policy evaluation estimators of the expected return in the inventory management domain based on 50 independent runs for each number of episodes. Normalisation is based on division by the squared range. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each number of episodes.
Episodes WIS WSIS WPDIS WSPDIS WINCRIS WSINCRIS SDRE SSDRE WDR WDRSIS
100 0.0490 0.0088 0.0278 0.0074 0.0641 0.0109 0.0160 0.0077 0.0173 0.0138
250 0.0651 0.0080 0.0335 0.0080 0.0546 0.0118 0.0161 0.0070 0.0132 0.0137
500 0.0652 0.0113 0.0340 0.0105 0.0456 0.0124 0.0163 0.0067 0.0098 0.0114
1000 0.0820 0.0122 0.0293 0.0119 0.0313 0.0128 0.0162 0.0066 0.0117 0.0104
TABLE IV: Mean squared error (MSE) for off-policy estimators of the expected return in the taxi domain based on 20 independent runs for each effective horizon. The best, second best, and third best estimators are highlighted in bold double-underlined, bold underlined, and bold, respectively, for each effective horizon.
Effective horizon WIS WSIS WPDIS WSPDIS WINCRIS WSINCRIS SDRE SSDRE WDR WDRSIS
10 0.6080 0.4525 0.7234 0.5084 0.4896 0.4767 1.2764 1.2790 53.2003 53.1011
50 15.8227 15.9281 11.6239 10.7777 35.6350 35.8282 9.4938 9.4435 188.8620 187.3876
250 440.6435 473.4986 281.9518 273.3647 1753.7938 1752.6067 25.3142 25.2920 993.3100 1055.8603
1000 8233.7773 7948.8306 4876.1759 4652.7797 27675.6630 27635.4627 231.6101 231.4313 11982.1668 12176.5372
VII-A Lift domains
Based on the $\epsilon$-negligibility definition, one would expect state-based techniques to be advantageous on domains where there are some states for which the action taken has no or limited impact on the return. This intuition is captured by the notion of “lift states”.
@@ VII-A Lift domains
Lift states. A lift state $s\in\mathcal{S}$ is a state for which $\mathcal{T}(s,a)=\mathcal{T}(s,a^{\prime})$ and $r(s,a)=r(s,a^{\prime})$ for all $a,a^{\prime}\in\mathcal{A}$. The set of lift states for a domain is denoted by $\mathcal{S}^{L}$.
Deterministic lift domain